# AutoGraphAD: A novel approach using Variational Graph Autoencoders for anomalous network flow detection

### v1

@@ 4.6 Anomaly Detection
The equation works by using a user-assigned score to weight the importance of each loss in the anomaly score. This allows us to use different weights for different settings. Additionally, we employ Robust Scaling [https://doi.org/10.1002/widm.1236] of all reconstruction errors and KL Divergence that are used for the anomaly scores. This guarantees that all scores are on the same scale and that no value dominates the range and skews the anomaly scores [10.1145/2594473.2594476]. Our entire approach runs in PyTorch and on the GPU. Providing a faster processing time that using Sci-kit learn [scikit-learn]. The pipeline can be found in Figure 9 in the Appendix.
4.7 Evaluating Inference Performance
To evaluate the performance of each method, we have chosen to measure the training and inference times. For timing, we used the Python built-in library, time. More specifically, we used the function perf_counter_ns() which gives us the most accurate measurement for our experiment, since we are measuring short durations.
For timing, we do not count the parts where data transfer takes place to make the comparisons more equal. Moreover, in the case of AutoGraphAD, we can time each individual pass when generating the Anomaly Score, but this is not possible during threshold setting and threshold inference timing and when timing Anomal-E’s downstream estimator training and inference. For these cases, we get the total time and divide it based on the number of graphs that were processed.
For the recorded timings, in the case of AutoGraphAD we add the average time per graph during the anomaly score with its respective time in the threshold component. For example, in the case of the training time, we add the average anomaly score generation time per graph with the estimated threshold calculation time per graph.
4.8 Libraries Used
To create our VGAE, we used PyTorch [NEURIPS2019_bdbca288] and an extension of PyTorch called PyTorch Geometric [fey2019fastgraphrepresentationlearning] which is best suited for the creation and training of GNNs. To ensure easier reproducibility and faster training, we employed PyTorch Lightning [Falcon_PyTorch_Lightning_2019]. Regarding anomaly detection, we employed NumPy [harris2020array] and PyOD [zhao2019pyod]. Finally, for data pre-processing, window creation, and running the downstream tasks for Anomal-E, we used Scikit-learn [scikit-learn] and Pandas [mckinney-proc-scipy-2010]. To run the Anomal-E code we used DGL [wang2020deepgraphlibrarygraphcentric].
5 Experimental Results
5.1 Experimental setup
To evaluate AutoGraphAD, we used UNSW-NB15 in both AutoGraphAD and Anomal-E. For Anomal-E, we used an implementation found in a GitHub repository [^1] https://github.com/waimorris/Anomal-E.
Overall, we trained the models in the different contamination levels [^2] The contamination level indicates the percentage of anomalous flows present in the training set. Ideally, anomaly detection methods should be trained only on normal flows. However, this is difficult to guarantee in real-world environments. Hence, robustness to a small degree of contamination is desirable for practical deployment in operational networks. followed by the anomaly detection estimators’ grid search for Anomal-E and weight search for the Anomaly Score in the VGAE implementation. During the hyperparameter search, the model weights were frozen.
@@ 5.1 Experimental setup
5.2 VGAE Variants
We exhaustively tested multiple different VGAE variants. Our base architecture is node masking and edge dropping to force the variational autoencoders to learn non-trivial node relationships. The variants that we evaluated are the following: $(1)$ Base architecture using MSE for node feature reconstruction, $(2)$ Base architecture using Cosine Embedding Error for node feature reconstruction, $(3)$ Regularised Base Architecture, $(4)$ Split Base Architecture, and $(5)$ Multiple Sampling Base Architecture.
On top of all of these architectures, we additionally evaluated each version using 1 and 2 layer encoders and decoders, as well as using KL annealing.
@@ 5.3 Training Hyperparameters
For AutoGraphAD, we have run multiple versions of our architecture. We have run different versions with different upgrades to evaluate which yields the best results. A summary of the different settings we used can be found in Table I.
Hyperparameter Hyperparameter Values
No. Layers [1, 2]
No. Hidden 32
Learning Rate 1e−3
Activation Function ReLU
Loss Function for Structural Loss Binary Cross Entropy
Loss Function for Feature Loss [Mean Square Error, Cosine Embedding Loss]
Optimiser AdamW
Weight Decay 1e-5
KL Annealing [Yes, No]
KL Annealing Epochs 10
KL Min Weight 0.0
Regulariser [Yes, No]
Multiple Draw [Yes, No]
Maximum Epochs 100
Early Stop Patience 20
Feature Importance 1.0
Structural Importance 1.0
TABLE I: AutoGraphAD Hyperparameters
5.4 Anomaly Detection Hyperparameters
For Anomal-E to work for anomaly detection, it requires a downstream detector to characterise attacks as anomalies or not. In the case of Anomal-E we chose to use the same algorithms as they used in their original paper [Caville_2022]. These algorithms are PCA, CBLOF, and HBOS. For the Hyperparameter search, we used a more extensive search than what was used in the original paper; a summary of the Hyperparameter search can be seen in Table II.
Hyperparameter Hyperparameter Values
PCA no. components [0.96, 0.98, 0.99]
PCA Weighted True
PCA Whiten [True, False]
PCA Standardisation True
HBOS Bins [5, 6, 8, 10, 12, 14, 16, 18, 20]
HBOS Alpha [0.05, 0.1]
HBOS Tol [0.1, 0.5]
Contamination (Anomal-E) [0.02, 0.035, 0.05, 0.1, 0.2]
Alpha (AutoGraphAD) [0.1, 0.5, 1.0]
Beta (AutoGraphAD) [0.1, 0.5, 1.0]
Gamma (AutoGraphAD) [0.1, 0.5, 1.0]
MSE use (AutoGraphAD) [True, False]
Percentile (AutoGraphAD) [95, 97, 98, 99]
TABLE II: Anomal-E Estimator and AutoGraphAD Anomaly Score Hyperparameters
For AutoGraphAD, we have tried different settings to find the combination that yields the most reliable anomaly scores. As mentioned in Section 4, we have individually configurable weights that affect the anomaly score. A detailed summary of the different parameters that we have evaluated can be found in Table II.
@@ 5.4 Anomaly Detection Hyperparameters
TABLE III: Results of Anomal-E and AutoGraphAD
@@
TABLE IV: Time measurement of the Anomal-E’s downstream estimators and AutoGraphAD. All times are the average from all the contamination levels and are the time per graph.
Estimator Training (s) Inference (s) Anomaly Score calculation (s) Threshold calculation (s) Threshold inference (s)
PCA [pyod_pca_module] 0.0885 0.0286
CBLOF [he2003cblof] 0.0454 0.0323
HBOS [goldstein2012hbos] 0.1359 0.0869
AutoGraphAD 0.0060 0.0046 0.00529 0.000015 0.000002
5.5 Results and interpretation
For the results, we have chosen the best of our variants for each level of contamination and different weights for the anomaly score. The model hyperparameters, anomaly score weights can be found in Tables VII and VIII in the Appendix. The metrics used are Accuracy, F1 Macro and Recall Macro. These metrics were chosen as an accurate descriptor of the model’s performance and because Macro treats each class as equal [du2023skewsensitiveevaluationframeworkimbalanced], whereas other types of results do not, making them ideal for evaluating the model in a highly unbalanced dataset like UNSW-Nb15 [7348942].
@@ 5.5 Results and interpretation
Figures 3, 4, and 5 show the results of our experimental evaluation for the different levels of contamination. AutoGraphAD yields very good results, comparable to those of Anomal-E. It should be noted that AutoGraphAD does not rely on any downstream anomaly detection algorithms such as Anomal-E, making it ideal for real-life deployment, as it can be tuned and adjusted more easily rather than requiring training on the new dataset. Furthermore, it should be noted that Anomal-E uses embeddings of size 256 [Caville_2022] whereas we rely on embeddings of size 32, 32 for the mean, and 32 for the Log Variance.
@@
Our results at 0% contamination exceed Anomal-E’s, especially when it comes to accuracy and F1 results. Anomal-E seems to be performing better in Recall, but looking at the F1 score its Precision seems to be lower than AutoGraphAD, indicating a higher false positive count than our own approach. This continues in the case of the 3.36% contamination level, where only CBLOF seems to perform better than AutoGraphAD. In the case of 5.76% contamination levels, again we achieve higher results with the exception of Recall against most of Anomal-E’s estimators with the exception of HBOS.
It is worth noting that the size of our embeddings is eight times smaller than those used by Anomal-E. This reduction is especially important for methods that use a downstream detector, because smaller embeddings can significantly improve both performance and processing speeds [doi:10.1126/science.153.3731.34].
Figure 6: Timing for inference and training in Seconds at 0% contamination in the training dataset.
Figure 7: The timings for training and inference at 3.36% contamination.
@@
AutoGraphAD yields substantially better inference time as seen in Table IV and Figures 6, 7, and 8. This can mainly be attributed to the fact that AutoGraphAD is GPU native, as it does not require a downstream detector as Anonmal-E, allowing for faster parallelisation and less time and resources spent on memory transfers to the CPU Anomal-E’s estimators.
Figure 8: The timings for training and inference at 5.76% contamination.
@@
Furthermore, AutoGraphAD is capable of being more easily tuned by just adjusting the weight parameters of the anomaly score, making alterations on the anomaly score generation on the fly. Unfortunately, this cannot truly be done in the case of the estimators as their characteristics are set during initialisation. Some of the estimator’ thresholds may be manually adjusted [zhao2019pyod] [scikit-learn], but this is also the case in AutoGraphAD. Finally, in the case of model retuning or retraining, Anomal-E’s estimators need to be re-trained from the start as the PyOD’s architecture does not support retuning [zhao2019pyod]. Instead, in our case, it is very easy to retune both the model and the anomaly score, since attacks are determined from a percentile threshold that is relatively lightweight in terms of inference and training. It should be noted that the use of Robust Scaling for each sliding window’s errors allows the model to be more resilient to anomalies, since extreme values do not dominate scaling [https://doi.org/10.1002/widm.1236]. Finally, traditional anomaly detection methods are not the best when it comes to streaming environments such as NIDS [app13106353], whereas AutoGraphAD scales each graph on its own and its thresholds and anomaly score weights can be easily retuned on the fly. This makes it an ideal candidate for an NIDS system.
6 Conclusion
@@
In conclusion, AutoGraphAD leverages recent advancements in the space of GVAEs and GAEs [GraphMAE, tian2023heterogeneousgraphmaskedautoencoders, kipf2016variationalgraphautoencoders, pan2019adversariallyregularizedgraphautoencoder] to build a more practical approach for network anomaly detection. Our design is relatively simple and yields results comparable to those of the Anomal-E. The main advantage of AutoGraphAD is faster inference and detection times, lower number of false positives, and easier and faster pipeline retuning than Anomal-E’s approach.
Furthermore, our evaluation allowed us to further explore the importance of node masking, negative edge sampling, and edge dropping. Our results highlight that lower masking, edge sampling, and edge dropping work better when we have very little or substantial contamination on our training graphs, whereas they seem more important when contamination is the same as with the rest of the dataset splits. This provides us with more information on how to tune our models for different training and inference conditions.
@@ Acknowledgements
This work was supported by Grant PCI2023-145974-2 funded by MICIU/AEI/10.13039/501100011033 and co-funded by the European Union (GRAPHS4SEC project).
@@
GitHub Copilot was used to help debug and develop the code. It served mainly as a search engine for what was causing the bugs and as a recommendation engine to check which libraries and their methods are the most suitable and optimal during development.
TABLE V: NetFlow Dataset Feature Set
# Column name Feature Value NF (v9)
1 src ip Source IP address Str Yes
2 src port Source Port Int Yes
3 dst ip Destination IP address Str Yes
4 protocol Destination port Int Yes
5 ip version IP version Int Yes
6 bidirectional first seen ms Timestamp in milliseconds of the first seen packet of the flow Int Yes
7 bidirectional last seen ms Timestamp in milliseconds of the last seen packet of the flow Int Yes
8 bidirectional duration ms Duration of the entire flow in milliseconds Int Yes
9 bidirectional packets Total packets of the flow Int Yes
10 bidirectional bytes Total bytes of the flow Int Yes
11 bidirectional min packet size Minimum packet size of the flow Int Yes
12 bidirectional max packet size Maximum packet size of the flow Int Yes
13 bidirectional mean packet size Mean of the size of the packet flows Float Yes
14 bidirectional mean packet iat ms Mean of the inter arrival packet time of all the packets of the flow in milliseconds Float Yes
15 bidirectional cumulative flags Logical OR of the flags of all the packets of the flow Int Yes
16 src2dst first seen ms Timestamp in milliseconds of the first seen packet from source to destination IP address Int Yes
17 src2dst last seen ms Timestamp in milliseconds of the last seen packet from source to destination IP address Int Yes
18 src2dst duration ms Total duration of the source to destination packets in milliseconds Int Yes
19 src2dst packets Total packets sent from source to destination IP address Int Yes
20 src2dst bytes Total bytes sent from source to destination IP address Int Yes
21 src2dst min packet size Minimum packet size sent from source to destination IP address Int Yes
22 src2dst max packet size Maximum packet size sent from source to destination IP address Int Yes
23 src2dst mean packet size Mean of the size of the packets sent from source to destination IP address Float Yes
24 src2dst mean packet iat ms Mean of the inter arrival packet time of all the packets sent from source to destination IP address in milliseconds Float Yes
25 src2dst cumulative flags Logical OR of the flags of all the packets sent from source to destination IP address Int Yes
26 dst2src first seen ms Timestamp in milliseconds of the first seen packet from destination to source IP address Int Yes
27 dst2src last seen ms Timestamp in milliseconds of the last seen packet from destination to source IP address Int Yes
28 dst2src duration ms Total duration of the destination to source packets in milliseconds Int Yes
29 dst2src packets Total packets sent from destination to source IP address Int Yes
30 dst2src bytes Total bytes sent from destination to source IP address Int Yes
31 dst2src min packet size Minimum packet size sent from destination to source IP address Int Yes
32 dst2src max packet size Maximum packet size sent from destination to source IP address Int Yes
33 dst2src mean packet size Mean of the size of the packets sent from destination to source IP address Float Yes
34 dst2src mean packet iat ms Mean of the inter arrival packet time of all the packets sent from destination to source IP address in milliseconds Float Yes
35 dst2src cumulative flags Logical OR of the flags of all the packets sent from destination to source IP address Int Yes
36 classification Labeling of the dataset (0 or 1) Int No
37 category Category of the classification if present Str No
TABLE VI: Anomal-E Estimator Model Hyperparameters
Estimator Contamination Level Hyperparameters
PCA No Contamination Contamination: 0.02, Number of Components: 0.96, Whiten: False
PCA 3.5% Contamination Contamination: 0.05, Number of Components: 0.98, Whiten: False
PCA 5.7% Contamination Contamination: 0.1, Number of Components: 0.98, Whiten: False
CBLOF No Contamination Contamination: 0.02, Alpha: 0.9, Beta: 5, Number of Clusters: 36, Use Weights: True
CBLOF 3.5% Contamination Contamination: 0.05, Alpha: 0.9, Beta: 5, Number of Clusters: 40, Use Weights: True
CBLOF 5.7% Contamination Contamination: 0.1, Alpha: 0.9, Beta: 5, Number of Clusters: 50, Use Weights: False
HBOS No Contamination Contamination: 0.02, Alpha: 0.05, Beta: 5, Number of Bins: 14, Tol: 0.1
HBOS 3.5% Contamination Contamination: 0.05, Alpha: 0.1, Beta: 5, Number of Bins: 5, Tol: 0.1
HBOS 5.7% Contamination Contamination: 0.1, Alpha: 0.1, Beta: 5, Number of Bins: 12, Tol: 0.1
Figure 9: Our Anomaly Score and classification pipeline. The pipeline very closes resembles the training of our model until the first part of the Anomaly Score generation, as once we have the losses, we scale them and combine them to generate the Anomaly Scores of each connection node. Once we have each node score, we pass them through a quantile search to calculate the necessary threshold that will from then on be used to determine if a connection is anomalous or not.
TABLE VII: VGAE Model Hyperparameters
Model Variant Contamination Model Hyperparameters
VGAE Regulariser 0% 20% Negative Sampling, Edge Dropping, Node Masking with Annealing, 1 Layer
VGAE Regulariser 3.5% 40% Negative Sampling, Edge Dropping, Node Masking with Annealing, 2 Layers
VGAE Multiple Draw 5.7% 20% Negative Sampling, Edge Dropping, Node Masking with Annealing, 1 Layer, 10 Draws
TABLE VIII: VGAE Anomaly Score Hyperparameters
Contamination Anomaly Score Hyperparameters
0% Contamination Alpha: 0.1, Beta: 0.5, Gamma: 0.1, MSE: True, Percentile: 95
3.5% Contamination Alpha: 1.0, Beta: 0.1, Gamma: 1.0, MSE: False, Percentile: 95
5.7% Contamination Alpha: 0.5, Beta: 0.1, Gamma: 0.5, MSE: True, Percentile: 95

### v2

@@ III-C Anomaly Detection
The equation works by using user-assigned weights to weigh the importance of each loss in the anomaly score. This allows us to use different weights for different settings and adapt to the changing nature of the network. We employ Robust Scaling [https://doi.org/10.1002/widm.1236] of all reconstruction errors and KL Divergence that are used for anomaly scores to ensure that no value dominates the scale. Our entire approach runs in PyTorch and on the GPU.
III-D Evaluating Inference Performance
To evaluate the performance of each method, we have chosen to measure the training and inference times. For timing, we used the Python built-in library, time.
We do not count the parts where data transfer takes place to make the comparisons more equal. Moreover, in the case of AutoGraphAD, we can time each individual pass by generating the Anomaly Score. However, we cannot time each pass in the threshold setting, threshold inference timing, and Anomal-E’s downstream estimator training and inference timing. In that form, we get the average pass time by dividing the total by all of the passes.
III-E Libraries Used
For AutoGraphAD, we used PyTorch, PyTorch Geometric [fey2019fastgraphrepresentationlearning] and DGL [wang2020deepgraphlibrarygraphcentric] for the GNNs. For faster training, we employed PyTorch Lightning [Falcon_PyTorch_Lightning_2019]. NumPy and PyOD [zhao2019pyod] are used for anomaly detection. For data pre-processing, window creation, and running downstream tasks for Anomal-E, we used Scikit-learn [scikit-learn] and Pandas.
IV Experimental Results
IV-A Experimental setup
To evaluate AutoGraphAD, we used UNSW-NB15 in both AutoGraphAD and Anomal-E. For Anomal-E, we used an implementation found in a GitHub repository [^1] https://github.com/waimorris/Anomal-E. We chose to only evaluate Anomal-E as it is a pure Anomaly Detection paper.
Overall, we trained the models in the different contamination levels followed by the anomaly detection estimators’ grid search for Anomal-E and weight search for the Anomaly Score in the VGAE implementation. During the hyperparameter search, the model weights were frozen. Labels are used only for offline evaluation and score-weight selection for anomaly detection.
@@ IV-A Experimental setup
IV-B VGAE Variants
We exhaustively tested multiple different VGAE variants. Our base architecture is node masking and edge dropping to force the variational autoencoders to learn non-trivial node relationships. The variants that we evaluated are the following: $(1)$ Base architecture using MSE for node feature reconstruction, $(2)$ Base architecture using Cosine Embedding Error for node feature reconstruction, $(3)$ Regularised Base Architecture, $(4)$ Split Base Architecture and $(5)$ Multiple Sampling Base Architecture.
We evaluated each version using 1 and 2 layer encoders and decoders, as well as using KL annealing. The implementation of all our variants are available in the accompanying code repository.
@@ IV-C Training Hyperparameters
For AutoGraphAD, we have run multiple versions of our architecture. We have run different versions with different upgrades to evaluate which yields the best results. We relied on Edge Index accuracy to select the best version of our model.
IV-D Anomaly Detection Hyperparameters
Anomal-E requires a downstream detector to characterise attacks as anomalies or not. In the case of Anomal-E we chose to use the same algorithms as they used in their original paper [Caville_2022]. For the Hyperparameter search, we used a more extensive search space, the grid can be found in our repository [^2] https://github.com/georgeani/AutoGraphAD/. Our Anomal-E implementation, including all variants evaluated, is available in our code repository. We used a Brute Force Hyperparameter search where we used the dataset labels to tune the parameters. We chose the results that yielded the best balance between F1-Score, Precision, and Recall.
@@ IV-D Anomaly Detection Hyperparameters
TABLE I: Results of Anomal-E and AutoGraphAD
For AutoGraphAD, we have tried different settings to find the combination that yields the most reliable anomaly scores. A detailed summary of the different parameters that we have evaluated can be found in our repository.
@@
IV-E Results and interpretation
For the results, we have chosen the best of our variants for each level of contamination and different weights for the anomaly score. The metrics used are Accuracy, F1 Macro and Recall Macro. These metrics were chosen as an accurate descriptor of the model’s performance and because Macro treats each class as equal [du2023skewsensitiveevaluationframeworkimbalanced], making them ideal for evaluating the model in a highly unbalanced dataset like UNSW-NB15 [7348942]. At 0% and 3.36% contamination, we used our regularised variant, and at 5.76% we used the multiple sampling variant.
@@ IV-E Results and interpretation
Figures 2, 3, and 4 show the results of our experimental evaluation for the different levels of contamination. AutoGraphAD yields very good results, comparable to those of Anomal-E. It should be noted that AutoGraphAD does not rely on any downstream anomaly detection algorithms such as Anomal-E, making it ideal for real-life deployment, as it can be tuned and adjusted more easily rather than requiring training on the new dataset.
@@
Our results at 0% contamination exceed Anomal-E’s, especially when it comes to accuracy and F1 results. Anomal-E seems to be performing better in Recall, but looking at the F1 score its Precision seems to be lower than AutoGraphAD, indicating a higher false positive count than our own approach. This continues at the 3.36% contamination level, where only CBLOF seems to perform better than AutoGraphAD. In the case of 5.76% contamination levels, again we achieve higher results with the exception of Recall against most of Anomal-E’s estimators.
@@
AutoGraphAD yields substantially better inference time (0.009892 s) compared to PCA (0.0286 s), CBLOF (0.0323 s) and HBOS(0.0869 s). This can also be observed in training: AutoGraphAD (0.011305 s), PCA (0.0885 s), CBLOF(0.0454 s), HBOS (0.1359 s). This can mainly be attributed to the fact that AutoGraphAD does not require a downstream detector as Anomal-E.
@@
AutoGraphAD is capable of being more easily tuned by simply adjusting the weight parameters of the anomaly score, making alterations to the anomaly score generation on the fly. Unfortunately, this cannot truly be done in the case of the estimators as their characteristics are set during initialisation. Some of the estimator’s thresholds may be manually adjusted [zhao2019pyod] [scikit-learn], but this is also the case in AutoGraphAD. Finally, in the case of model retuning or retraining, Anomal-E’s estimators need to be re-trained from the start as the PyOD’s architecture does not support retuning [zhao2019pyod]. Instead, AutoGraphAD is very easy to retune both the model and the anomaly score due to its design.
V Conclusion
@@
In conclusion, AutoGraphAD leverages recent advancements in the space of GVAEs and GAEs [GraphMAE, tian2023heterogeneousgraphmaskedautoencoders, kipf2016variationalgraphautoencoders, pan2019adversariallyregularizedgraphautoencoder] to build a more practical approach for network anomaly detection. Our design is relatively simple and yields results comparable to those of the Anomal-E. The main advantage of AutoGraphAD is faster inference and detection times, a lower number of false positives, and easier and faster pipeline retuning than Anomal-E’s approach.
@@ Acknowledgements
This work was supported by Grant PCI2023-145974-2 funded by MICIU/AEI/10.13039/501100011033 and cofunded by the European Union (GRAPHS4SEC project). This work is also supported by the Catalan Institution for Research and Advanced Studies (ICREA Academia). GitHub Copilot was used to help debug and develop the code. It served mainly as a search engine for what was causing the bugs and as a recommendation engine to check which libraries and their methods are the most suitable and optimal during development.